OpenStudy (anonymous):

PLEASE HELP I WILL MEDAL AND FAN!!! A large smartphone company produces 15,000 smartphones per hour. How long would it take this company to produce 300,000 smartphones? A. 15 h B. 20 h C. 30 h D. 40 h

OpenStudy (anonymous):

@welshfella

OpenStudy (texaschic101):

15000/1 = 300000/x --- 15000 per hr = 300000 per x hrs cross multiply and solve for x

OpenStudy (anonymous):

Ok can you make that simpler to understand because this is hard for me to understand.

OpenStudy (anonymous):

B.20

OpenStudy (anonymous):

Thank you @Ortegafrancisco but I need to know how to work it out if you just tell me it wont help me with future questions like this...My point is I don't want direct answers I want somebody to help me understand how to work it out. Ok?

OpenStudy (texaschic101):

 sorry...I can't draw very good on this I just cross multiplied....now x = 300000/15000 which equals what ?